OpenStudy (anonymous):

Help I am stuck on like 4 of my geometry study guide questions! :( I will medal and fan for help .

OpenStudy (anonymous):

Can we see the questions? Thanks. XP

OpenStudy (anonymous):

yes questions

OpenStudy (anonymous):

okay here is one of the questions.

OpenStudy (anonymous):

sorry my computer is really slow so it takes me a lil bit longer to respond.

OpenStudy (anonymous):

>??????????

OpenStudy (anonymous):

@superbatman2

Directrix (directrix):

Because the two rectangles are similar, the ratio of corresponding sides is the same. That means this: 58.5 is to 13 as x is to 16. x is the longer side of the larger rectangle.

Directrix (directrix):

58.5 / 13 = x / 16 Solve the proportion for x.

Directrix (directrix):

x is not the answer to the question. Post what you get for the value of x and then we will continue. Hint: Cross multiply to begin to find x.

OpenStudy (anonymous):

13x/936 ?

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

13*x = 16 * 58.5

OpenStudy (anonymous):

x=72?

Directrix (directrix):

x = 72 http://www.wolframalpha.com/input/?i=13*x+%3D+16+*+58.5

Directrix (directrix):

There are 5 strips of molding. 5* 36 = ?

OpenStudy (anonymous):

ok

Directrix (directrix):

5* 36 = does not equal 72. Try again.

OpenStudy (anonymous):

180

Directrix (directrix):

@kyley_bear Need your help. Let's crank this out and be finished. Please answer the question.

Directrix (directrix):

Tag me when you do the multiplication.

Directrix (directrix):

So, take the sum of the perimeters of the two rectangles and subtract the 180 from that: 2 * 58.5 + 2*72 + 2*16 + 2*13 = perimeter of the two rectangles

OpenStudy (anonymous):

@Directrix

Directrix (directrix):

Required molding to finish the framing: (2 * 58.5 + 2*72 + 2*16 + 2*13) - 180 =

Directrix (directrix):

That is 139 but do check. I get that your computer is lagging. http://www.wolframalpha.com/input/?i=+%282+*+58.5+%2B+2*72+%2B+2*16+%2B+2*13%29+-+180+%3D

Directrix (directrix):

Post the equation you used to get the answer you just deleted, and I'll check it against the one I had before this site crashes again.

OpenStudy (anonymous):

I redid it and I got your answer I missed 2*16 thats why I got 107 lol

Directrix (directrix):

Okay. Cool that you self-diagnosed the error.